OpenStudy (anonymous):

find the vertex form of the following equation by completing the square. y=3x^2+6x+8 A.y=(x+4)^2 B.y=3(x+2)^2+4 C. y=3(x+1)^2+5 D. y=(x+3)^2+5

OpenStudy (jdoe0001):

why not just roll a dice?

OpenStudy (dumbcow):

\[3x^{2} +6x = 3(x^{2}+2x) = 3(x^{2}+2x+1) -3 = 3(x+1)^{2} -3\]